Please show work and thanks in advance.
Find the equation of the parabola with the given focus and directrix. Focus (5,2), directrix y=3

Answer by lwsshak3(11628)

Find the equation of the parabola with the given focus and directrix. Focus (5,2), directrix y=3
This is a parabola that opens downwards.
Its standard form of equation:
(x-h)^2=-4p(y-k), (h,k) = (x,y) coordinates of the vertex
For given parabola:
vertex:(5,5/2)
axis of symmetry:x=5
p=1/2 (distance from directrix or focus to vertex on the axis of symmetry.
4p=2
Equation:
(x-5)^2=-2(y-5/2)